# Definability of band structures on posets

### v1

@@ Abstract
As an application of the order-theoretic perspective on bands, we generalize results by the third author, obtaining “inner” direct product representations for RRBs having a central (commuting) element.
@@ 1 Introduction
[…]
In this paper, we are interested in studying the class of associative posets, which comprises all partial orders underlying right-regular bands. In doing so, we will interpret some constructions that give rise to RRBs from an order-theoretical point of view. This vantage points proves to be useful in a number of situations, e.g. the analysis of direct product decompositions carried out in Section 4. Another example is given by the results in [3], which characterizes all varieties of bands according to some particular congruences.
@@
This discussion of definability permeates our whole work, and many questions remain open; some of these are gathered in Section 5. After setting up some preliminaries in the next section, we restrict ourselves in Section 2.2 to a subfamily of associative posets for which a neat structural characterization is available.
2 Associative posets
2.1 Preliminaries
@@
for all $x,y$ in $P$. In the case of an associative poset, we can moreover choose this to be an RRB operation and then we use the word posemigroup when referring to the expanded structure $(P,\leq,\cdot)$. Conversely, given an RRB $(P,\cdot)$, we say that the partial order $x\leq y$ given by (2) is the underlying order of this RRB.
We start by introducing some useful elementary properties of posemigroups.
@@
    From Item 1 we know that $a\cdot(b\cdot a)\leq b\cdot a$ and $b\cdot a=b\cdot(a\cdot b\cdot a)\leq a\cdot b\cdot a$. By antisimmetry we obtain $a\cdot b\cdot a=b\cdot a$.
@@
Every decreasing subset of a posemigroup is a substructure.
@@
Let $P$ be a posemigroup. If $x\cdot y=y$ then $(y\cdot x){\downarrow}$ is isomorphic to $y{\downarrow}$.
@@
We prove that the function $f:(y\cdot x){\downarrow}\to y{\downarrow}$ defined by $f(a)\mathrel{\mathop{:}}=a\cdot y$ is an isomorphism with inverse $b\mapsto b\cdot x$. The injectivity of $f$ follows from the fact that for all $a\leq y\cdot x$, we have $a=a\cdot(y\cdot x)=(a\cdot y)\cdot x$. Now, if $b\leq y$, take $a=b\cdot x$. We know $a\leq y\cdot x$ by Lemma 2.3(1). Now, $f(a)=(b\cdot x)\cdot y=b\cdot(x\cdot y)=b\cdot y=b$. So $f$ is surjective and hence a bijection. The fact that $f$ is order preserving is a direct consequence of Lemma 2.3(1). To see that $f^{-1}$ is order preserving, take $a\leq b\leq y$. By Lemma 2.3(1) $f^{-1}(a)=a\cdot x\leq b\cdot x=f^{-1}(b)$. So $f$ is an isomorphism. ∎
@@
Let $P$ be a posemigroup. For every $x,y\in P$, $(x\cdot y){\downarrow}$ is isomorphic to $(y\cdot x){\downarrow}$.
@@
Let $P$ be a posemigroup and $m\in P$ be minimal. For every $x\in P$, $m\cdot x\leq x$ is minimal.
2.2 Normal posets
@@
Our goal for this section is to give a characterization of normal posets, which are the orders underlying right-normal bands. Normal posets are always relative meet-semilattices (i.e., posets in which every principal ideal is a meet-semilattice) but not conversely (even assuming associativity; see Example 2.21). In fact, it is not hard to show that an associative relative meet-semilattice is a normal poset if and only if it admits an operation which acts as the meet operation in every principal ideal [4, Lemma 10].
First, we prove an auxiliary result.
@@
Let $A$ be an RRB. Then $\theta:=\{(x,y)\in A^{2}:x\cdot y=y\mathrel{\&}y\cdot x=x\}$ is a congruence over $A$ and $A/\theta$ is a meet-semilattice.
Proof.
The relation $\theta$ is clearly symmetric and reflexive. If $x\cdot y=y$ and $y\cdot z=z$, then $x\cdot z=x\cdot(y\cdot z)=(x\cdot y)\cdot z=y\cdot z=z$. Analogously $z\cdot x=x$. Thus $\theta$ is transitive. Suppose now that $\langle x,x^{\prime}\rangle,\langle y,y^{\prime}\rangle\in\theta$. Then $(x\cdot y)\cdot\underline{(x^{\prime}\cdot y^{\prime})}=(x\cdot y)\cdot(y^{\prime}\cdot x^{\prime}\cdot y^{\prime})=x\cdot\underline{y\cdot y^{\prime}}\cdot x^{\prime}\cdot y^{\prime}=x\cdot\underline{y^{\prime}\cdot x^{\prime}\cdot y^{\prime}}=\underline{x\cdot x^{\prime}}\cdot y^{\prime}=x^{\prime}\cdot y^{\prime}$. Therefore $\theta$ is a congruence. To see that $A/\theta$ is a meet-semilattice, note that for every $x,y\in A$, $(x\cdot y)\cdot(y\cdot x)=x\cdot\underline{y\cdot y}\cdot x=\underline{x\cdot y\cdot x}=y\cdot x$. Analogously, $(y\cdot x)\cdot(x\cdot y)=x\cdot y$, thus $\langle x\cdot y,y\cdot x\rangle\in\theta$. Therefore, we have that $x/\theta\cdot y/\theta=(x\cdot y)/\theta=(y\cdot x)/\theta=y/\theta\cdot x/\theta$. So $A/\theta$ is a meet-semilattice. ∎
In view of this result, for any RRB $A$, we shall call $\theta$ the semilattice congruence of $A$ and $A/\theta$ the quotient semilattice of $A$.
@@
($\Rightarrow$) Fix an admissible RNB operation $\cdot$ for $P$. Let $\theta$ be the semilattice congruence of $(P,\cdot)$ and $f:P\rightarrow P/\theta$ the canonical projection. First note that since each initial segment $p{\downarrow}$ is decreasing, it is also a substructure, so the restriction of any homomorphism to it is a homomorphism. We also have that RRB isomorphisms are isomorphisms of the underlying orders. Therefore we only need to check that $f|_{p{\downarrow}}=f_{p}$ is bijective for every $p\in P$. Let $p\in P$; since $(P,\cdot)$ is an RNB, if $x,y\leq p$ and $f(x)=f(y)$, then
@@
[…]
@@
$B$ is a subalgebra of the direct product RNB structure on $S\times A$. Let $h:B\rightarrow P$ given by $h(\langle x,m\rangle)=(f_{m})^{-1}(x)$. Note that $h$ is surjective.
Claim 1.
@@
Figure 2: Multiple RNB structures
[…]
$\varphi(x,y,z):=z\leq y\wedge\psi(z,x)\wedge\forall d,\;(\psi(d,x)\wedge\psi(d,y))\rightarrow\psi(d,z).$
@@
A few questions regarding definability remain open. One of them is the following:
Question 2.12.
2.3 Constructions
@@
First, we prove that the sum of two associative posets $P$ and $Q$ is associative. We define a posemigroup operation over $P+Q$ in the following manner:
@@
Let $x,y,z$ be arbitrary elements in $\bigsqcup_{i\in I}P_{i}$ and $i,j,k\in I$ be such that $x\in P_{i}$, $y\in P_{j}$, $z\in P_{k}$. Now we have, by associativity of $P_{i}+P_{j}+P_{k}$ and the fact that this finite sum is a subalgebra of $\sum_{i\in I}P_{i}$, that $(x\cdot y)\cdot z=x\cdot(y\cdot z)$. ∎
@@ 2.4 Examples
The poset in Figure 4 is not associative. Assume by way of contradiction that it admits an RRB structure $\cdot$. By Lemma 2.3(2), $b\cdot x=x$ and $x\cdot b=b$; but this contradicts Corollary 2.5.
@@ 2.5 Foliated trees
1$\Rightarrow$2 and 2$\Rightarrow$3 are trivial. Let us prove 3$\Rightarrow$4. Let $\mathcal{F}$ be a non empty family of mutually disjoint non empty sets. We now define a tree order over ${}T\mathrel{\mathop{:}}=\{\mathcal{F}\}\cup\mathcal{F}\cup\bigcup\mathcal{F}$: $x<y$ if and only if $x\neq y$, and $y=\mathcal{F}$ or $x\in y$ (we are considering $\mathcal{F}\cap\bigcup\mathcal{F}=\varnothing$. If this was not the case, the order obtained would not be that of a tree with height $3$. This can be fixed by considering {}×$\{\mathcal{F}\}\times\{2\}\cup\mathcal{F}\times\{1\}\cup\mathop{\textstyle\bigcup}\mathcal{F}\times\{0\}$ as the universe for $T$ and defining the order as: $x^{\prime}=\langle x,n\rangle<y^{\prime}=\langle y,m\rangle$ if and only if $y^{\prime}=\langle\mathcal{F},2\rangle$ or $x\in y$ and $n<m$). Note that this is a tree with height $3$ and therefore it is associative. Fix an admissible RRB structure for $T$ and a minimal $m\in T$. Then $\{m\cdot B:B\in\mathcal{F}\}$ is transversal for $\mathcal{F}$ as for $B\in\mathcal{F}$, $m\cdot B$ is a minimal element below $B$ by Corollary 2.6.
@@
    This classification is exhaustive because by  2.24 $z$ and $y$ must be comparable.
@@
      In this case we know by 2.24 that $x$ and $z$ must be incomparable as $x$ and $y$ are so: $x\leq z$ would imply $x\leq y$, as $z\leq x$, together with  2.24 would imply that $x$ and $y$ are comparable. We also have that $F(y)=F(z)$ by Claim 4. Then $x\cdot(y\cdot z)=x\cdot z=F(z)=F(y)=F(y)\cdot z=(x\cdot y)\cdot z$.
@@
The previous theorem still holds if we chose $T$ to be a forest of foliated trees instead of just a foliated tree. This result follows from the fact that if we have a forest of foliated trees $T$, then we can consider the poset $T^{\prime}:=T+\{1\}$, where $1\notin T$. $T^{\prime}$ is a foliated tree and is therefore associative by the previous theorem. Finally, since $T$ is a decreasing subset of $T^{\prime}$, it is associative; in fact, it is a subalgebra with respect to the operation we defined in the previous theorem.
@@
Let $z$ be such a lower bound. That is $z\leq x$ and $z\leq y$. Then $f(z)\leq f(x)$ and $f(z)\leq f(y)$. As the codomain of $f$ is a forest, $f(x)$ and $f(y)$ are comparable by  2.24. If $f(x)\neq f(y)$, then either $f(x)<f(y)$ or $f(y)<f(x)$. In both of these cases $x$ and $y$ are comparable. ∎
Let’s take an element $1$ not belonging to $T$. Let $\tilde{F}:T+\{1\}\rightarrow T+\{1\}$ a function defined like the one in 2.25. That is, a function such that $\tilde{F}(a)$ is minimal below $a$ for ever $a$ in $T$, and also, for every minimal $b\in T$, $\tilde{F}^{-1}(b)$ is a convex chain containing $b$.
@@
If $f(y)=f(z)$, the result holds trivially. Otherwise, we have that $f(z)<f(y)$. Note that $f(F(y))=\tilde{F}(f(y))\leq f(z)<f(y)$ and therefore by 4, $\tilde{F}(f(z))=\tilde{F}(f(y))$. This tells us, by definition of $F$, that $F(z)=F(y)$. ∎
@@
Consider that $f(x)=f(y)\neq f(z)$ $x\leq y$ and $y,z$ incomparable. Then $x\cdot(y\cdot z)=x\cdot F(z)=F(z)=x\cdot z=(x\cdot y)\cdot z$ because $x$ and $z$ are incomparable. This is because, if $x\leq y$ and $y$ is incomparable with $z$, it cannot be the case that $z\leq x$. It also cannot be $x\leq z$ because that would imply that ${y,z}$ has a lower bound, which contradicts 6. Now consider that $f(x)\neq f(y)=f(z)$ $x,y$ incomparable,$y\leq z$. Then $(x\cdot y)\cdot z=F(y)\cdot z=F(y)=F(y\mathbin{\tilde{\cdot}}z)=x\cdot(y\mathbin{\tilde{\cdot}}z)=x\cdot(y\cdot z)$ because $F(y)\leq z$, $F(y)=F(y\mathbin{\tilde{\cdot}}z)$, and because $x$ is incomparable with $y\mathbin{\tilde{\cdot}}z$ (since it cannot be $x\leq y\mathbin{\tilde{\cdot}}z$ by  5, and it cannot be $y\mathbin{\tilde{\cdot}}z\leq x$ because that, together with  6, would imply $z\leq x$ and that would mean $y\leq x$). ∎

### v2

@@ 1 Introduction
The associated equivalence relation, $\mathfrak{L}\mathrel{\mathop{:}}=(\lesssim)\cap(\gtrsim)$ (corresponding to Green’s equivalence $\mathfrak{l}$ [3]) is not always a congruence relation, but only for “left-semiregular” ones [5]. In particular, the variety of right-regular bands (RRB), axiomatized by the equation $x\cdot y\cdot x=y\cdot x$, are characterized by the fact that $\mathfrak{l}$ is the identity and hence the quasiorder (1) is actually a partial order.
In this paper, we are interested in studying the class of associative posets, which comprises all partial orders underlying right-regular bands. In doing so, we will interpret some constructions that give rise to RRBs from an order-theoretical point of view. This vantage point proves to be useful in a number of situations, e.g. the characterization of all varieties of bands according to some particular congruences [5].
@@
This discussion of definability permeates our whole work, and many questions remain open; some of these are gathered in Section 7. After setting up some preliminaries in the next section, we restrict ourselves in Section 3 to a subfamily of associative posets for which a neat structural characterization is available.
2 Preliminaries
@@
for all $x,y$ in $P$. In the case of an associative poset, we can moreover choose this to be an RRB operation and then we use the phrase right posemigroup when referring to the expanded structure $(P,\leq,\cdot)$; this stems from the fact that the product is order-preserving on the right only (see Lemma 2.5.1). Conversely, given an RRB $(P,\cdot)$, we say that the partial order $x\leq y$ given by (2) is the underlying order of this RRB.
We start by introducing some useful elementary properties of right posemigroups.
@@
    From Item 1 we know that $a\cdot(b\cdot a)\leq b\cdot a$ and $b\cdot a=b\cdot(a\cdot b\cdot a)\leq a\cdot b\cdot a$. By antisymmetry we obtain $a\cdot b\cdot a=b\cdot a$.
@@
Definition 2.2.
Let $P$ be a poset and $D$ a subset of $P$. We say that $D$ is a decreasing subset of $P$ if $\forall a,b\in P,a\in D\mathrel{\&}b\leq a\Rightarrow b\in D$.
Definition 2.3.
Let $A$ be a right posemigroup and $B$ a subset of $A$. We say that $B$ is a substructure of $A$ if $B$ is closed under the semigroup operation.
@@
Every decreasing subset of a right posemigroup is closed under the semigroup operation, and hence it is the universe of a substructure.
@@
Let $P$ be a right posemigroup. If $x\cdot y=y$ then $(y\cdot x){\downarrow}$ is isomorphic to $y{\downarrow}$.
@@
We prove that the function $f:(y\cdot x){\downarrow}\to y{\downarrow}$ defined by $f(a)\mathrel{\mathop{:}}=a\cdot y$ is an isomorphism with inverse $b\mapsto b\cdot x$. The injectivity of $f$ follows from the fact that for all $a\leq y\cdot x$, we have $a=a\cdot(y\cdot x)=(a\cdot y)\cdot x$. Now, if $b\leq y$, take $a=b\cdot x$. We know $a\leq y\cdot x$ by Lemma 2.5(1). Now, $f(a)=(b\cdot x)\cdot y=b\cdot(x\cdot y)=b\cdot y=b$. So $f$ is surjective and hence a bijection. To see that $f$ preserves the right posemigroup operation, take $a,b\in(y\cdot x){\downarrow}$. We have that $f(a\cdot b)=(a\cdot b)\cdot y=a\cdot(b\cdot y)=a\cdot(y\cdot b\cdot y)=(a\cdot y)\cdot(b\cdot y)=f(a)\cdot f(b)$. ∎
@@
Let $P$ be a right posemigroup. For every $x,y\in P$, $(x\cdot y){\downarrow}$ is isomorphic to $(y\cdot x){\downarrow}$.
@@
Let $P$ be a right posemigroup and $m\in P$ be minimal. For every $x\in P$, $m\cdot x\leq x$ is minimal.
3 Normal posets
@@
Our goal for this section is to give a characterization of normal posets, which are the orders underlying right-normal bands. Normal posets are always relative meet-semilattices (i.e., posets in which every principal ideal is a meet-semilattice) but not conversely (even assuming associativity; see Example 5.4). In fact, it is not hard to show that an associative relative meet-semilattice is a normal poset if and only if it admits an operation which acts as the meet operation in every principal ideal [6, Lemma 10].
First, we require an auxiliary result. Recall that in every band $B$, Green’s equivalence $\mathfrak{R}$ is characterized by $x\cdot y=y\mathrel{\&}y\cdot x=x$. Also, the equivalence $\mathfrak{D}$ given by the composition of $\mathfrak{R}$ and $\mathfrak{L}$, is the least congruence such that $B/\mathfrak{D}$ is a semilattice [4]. Since on every right regular band, Green’s equivalence $\mathfrak{L}$ is trivial, we have $\mathfrak{D}$ = $\mathfrak{R}$ on right regular bands. From this, we can obtain the following Lemma:
@@
Let $B$ be an RRB. Then $\mathfrak{D}=\{(x,y)\in B^{2}:x\cdot y=y\mathrel{\&}y\cdot x=x\}$ is the least congruence such that $B/\mathfrak{D}$ is a semilattice.
In view of this result, for any RRB $B$, we shall call $\mathfrak{D}$ the semilattice congruence of $B$ and $B/\mathfrak{D}$ the quotient semilattice of $B$.
@@
($\Rightarrow$) Fix an admissible RNB operation $\cdot$ for $P$. Let $\mathfrak{D}$ be the semilattice congruence of $(P,\cdot)$ and $f:P\rightarrow P/\mathfrak{D}$ the canonical projection. First note that since each initial segment $p{\downarrow}$ is decreasing, it is also a substructure, so the restriction of any homomorphism to it is a homomorphism. We also have that RRB isomorphisms are isomorphisms of the underlying orders. Therefore we only need to check that $f|_{p{\downarrow}}=f_{p}$ is bijective for every $p\in P$. Let $p\in P$; since $(P,\cdot)$ is an RNB, if $x,y\leq p$ and $f(x)=f(y)$, then
@@
since $x\mathrel{\mathfrak{D}}y$ and hence $f_{p}$ is injective. Also, if $x/\mathfrak{D}\leq p/\mathfrak{D}$, then $x\cdot p/\mathfrak{D}=x/\mathfrak{D}$, so $f_{p}(x\cdot p)=x/\mathfrak{D}$. Thus $f_{p}$ is an order isomorphism for every $p\in P$.
@@
$B$ is a substructure of the direct product RNB structure on $S\times A$. Let $h:B\rightarrow P$ given by $h(\langle x,m\rangle)=(f_{m})^{-1}(x)$. Note that $h$ is surjective.
Claim 3.3.
@@
Proof.
Consider the equivalence relation over the previous poset in which the only non-trivial pair consists of the two maximal elements of the poset. The quotient of this relation is a meet-semilattice and the canonical projection satisfies the conditions of the previous theorem. ∎
@@
Figure 2: Multiple RNB structures over a normal poset. We define three different equivalence relations. Two elements are related if and only if their nodes are of the same color and shape. One can easily check that the quotient and the canonical projection satisfy the hypotheses of the previous theorem in each case.
Note that the quotient semilattice of the RNB obtained in the last part of the proof is isomorphic to the semilattice $S$ with which we started. Also the canonical projection of the semilattice congruence is the homomorphism $f$. It is not hard to see that in the case of right-normal bands, one can completely determine the product using two equationally definable binary relations: the underlying partial order and the semilattice congruence. This tells us, according to Beth’s Theorem, that there is a first order formula $\varphi(x,y,z)$ in the language $\{\leq,\mathfrak{D}\}$ which is equivalent to the formula $x\cdot y=z$. In fact, Theorem 3.2 provides information for defining it. We simply need a formula which roughly says “$z$ is the representative of the class $x/\mathfrak{D}\wedge y/\mathfrak{D}$ which is below $y$”. For this, we define $\psi(x,y)$ which states “$x/\mathfrak{D}\leq y/\mathfrak{D}$”, as $\psi(x,y):=\exists c,\;c\leq y\mathrel{\&}x\mathrel{\mathfrak{D}}c$. Then we set
$\varphi(x,y,z):=z\leq y\mathrel{\&}\psi(z,x)\mathrel{\&}\forall d,\;(\psi(d,x)\mathrel{\&}\psi(d,y))\rightarrow\psi(d,z).$
@@
It’s easy to see that the previous poset is not a normal poset as it is not even a relative meet-semilattice. Associativity stems from Remark 6.7.
4 Constructions
@@
First, we prove that the sum of two associative posets $P$ and $Q$ is associative. We define a right posemigroup operation over $P+Q$ in the following manner:
@@
Let $x,y,z$ be arbitrary elements in $\bigsqcup_{i\in I}P_{i}$ and $i,j,k\in I$ be such that $x\in P_{i}$, $y\in P_{j}$, $z\in P_{k}$. Now we have, by associativity of $P_{i}+P_{j}+P_{k}$ and the fact that this finite sum is a substructure of $\sum_{i\in I}P_{i}$, that $(x\cdot y)\cdot z=x\cdot(y\cdot z)$. ∎
@@ 5 Examples
The poset in Figure 4 is not associative. Assume by way of contradiction that it admits an RRB structure $\cdot$. By Lemmas 2.5(2) and 2.1(1), $b\cdot x=x$ and $x\cdot b=b$; but this contradicts Corollary 2.7, since $x{\downarrow}$ and $b{\downarrow}$ are obviously not isomorphic.
@@ 6 Foliated trees
1$\Rightarrow$2 and 2$\Rightarrow$3 are trivial. Let us prove 3$\Rightarrow$4. We will show, using 3, that every non empty family $\mathcal{F}$ of mutually disjoint non empty sets has a transversal. Let $\mathcal{F}$ be such a family. We now define a tree order over ${}T\mathrel{\mathop{:}}=\{\mathcal{F}\}\cup\mathcal{F}\cup\bigcup\mathcal{F}$: $x<y$ if and only if $x\neq y$, and $y=\mathcal{F}$ or $x\in y$ (we are considering $\mathcal{F}\cap\bigcup\mathcal{F}=\varnothing$. If this was not the case, the order obtained would not be that of a tree with height $3$. This can be fixed by considering {}×$\{\mathcal{F}\}\times\{2\}\cup\mathcal{F}\times\{1\}\cup\mathop{\textstyle\bigcup}\mathcal{F}\times\{0\}$ as the universe for $T$ and defining the order as: $x^{\prime}=\langle x,n\rangle<y^{\prime}=\langle y,m\rangle$ if and only if $y^{\prime}=\langle\mathcal{F},2\rangle$ or $x\in y$ and $n<m$). Note that this is a tree with height $3$ and therefore it is associative. Fix an admissible RRB structure for $T$ and a minimal $m\in T$. Then $\{m\cdot B:B\in\mathcal{F}\}$ is transversal for $\mathcal{F}$ as for $B\in\mathcal{F}$, $m\cdot B$ is a minimal element below $B$ by Corollary 2.8.
@@
    This classification is exhaustive because by Proposition 6.1 $z$ and $y$ must be comparable.
@@
      In this case we know by Proposition 6.1 that $x$ and $z$ must be incomparable as $x$ and $y$ are so: $x\leq z$ would imply $x\leq y$, as $z\leq x$, together with Proposition 6.1 would imply that $x$ and $y$ are comparable. We also have that $F(y)=F(z)$ by Claim 6.4. Then $x\cdot(y\cdot z)=x\cdot z=F(z)=F(y)=F(y)\cdot z=(x\cdot y)\cdot z$.
@@
The previous theorem still holds if we chose $T$ to be a forest of foliated trees instead of just a foliated tree. This result follows from the fact that if we have a forest of foliated trees $T$, then we can consider the poset $T^{\prime}:=T+\{1\}$, where $1\notin T$. $T^{\prime}$ is a foliated tree and is therefore associative by the previous theorem. Finally, since $T$ is a decreasing subset of $T^{\prime}$, it is associative; in fact, it is a substructure with respect to the operation we defined in the previous theorem.
@@
Let $z$ be such a lower bound. That is $z\leq x$ and $z\leq y$. Then $f(z)\leq f(x)$ and $f(z)\leq f(y)$. As the codomain of $f$ is a forest, $f(x)$ and $f(y)$ are comparable by Proposition 6.1. If $f(x)\neq f(y)$, then either $f(x)<f(y)$ or $f(y)<f(x)$. In both of these cases $x$ and $y$ are comparable. ∎
Let’s take an element $1$ not belonging to $T$. Let $\tilde{F}:T+\{1\}\rightarrow T+\{1\}$ a function defined like the one in Theorem 6.2. That is, a function such that $\tilde{F}(a)$ is minimal below $a$ for ever $a$ in $T$, and also, for every minimal $b\in T$, $\tilde{F}^{-1}(b)$ is a convex chain containing $b$.
@@
If $f(y)=f(z)$, the result holds trivially. Otherwise, we have that $f(z)<f(y)$. Note that $f(F(y))=\tilde{F}(f(y))\leq f(z)<f(y)$ and therefore by 6.4, $\tilde{F}(f(z))=\tilde{F}(f(y))$. This tells us, by definition of $F$, that $F(z)=F(y)$. ∎
@@
Consider that $f(x)=f(y)\neq f(z)$ $x\leq y$ and $y,z$ incomparable. Then $x\cdot(y\cdot z)=x\cdot F(z)=F(z)=x\cdot z=(x\cdot y)\cdot z$ because $x$ and $z$ are incomparable. This is because, if $x\leq y$ and $y$ is incomparable with $z$, it cannot be the case that $z\leq x$. It also cannot be $x\leq z$ because that would imply that ${y,z}$ has a lower bound, which contradicts 6.10. Now consider that $f(x)\neq f(y)=f(z)$ $x,y$ incomparable,$y\leq z$. Then $(x\cdot y)\cdot z=F(y)\cdot z=F(y)=F(y\mathbin{\tilde{\cdot}}z)=x\cdot(y\mathbin{\tilde{\cdot}}z)=x\cdot(y\cdot z)$ because $F(y)\leq z$, $F(y)=F(y\mathbin{\tilde{\cdot}}z)$, and because $x$ is incomparable with $y\mathbin{\tilde{\cdot}}z$ (since it cannot be $x\leq y\mathbin{\tilde{\cdot}}z$ by  6.9, and it cannot be $y\mathbin{\tilde{\cdot}}z\leq x$ because that, together with  6.10, would imply $z\leq x$ and that would mean $y\leq x$). ∎